OpenStudy (anonymous):

find f'(x)=5+e^x-e^x(x)/(5+e^x)^2

OpenStudy (sgadi):

please give equation clearly ...

OpenStudy (anonymous):

find the derivative of 5+e^x(e to the power of x) - e^x (x)/ Divided by (5+e^x)^2 (5plus e to power of 2)all sqaured

OpenStudy (sgadi):

A \[\frac{5+e^x-e^{x^{x}}}{\left( 5+e^x\right)^2}\] b \[\frac{5+e^x-e^{x^2}}{\left( 5+e^x\right)^2}\] which one is you are asking?

OpenStudy (anonymous):

A but its not e\[^{x}^{x} its (x)e ^{x}\]

OpenStudy (anonymous):

5+e^x-(x)e^x

OpenStudy (anonymous):

then divide that by (5+e^x)^2

OpenStudy (sgadi):

f=(5+exp(x)+x*exp(x))/((5+exp(x))^2) then df/dx=(2*exp(x) + x*exp(x))/(exp(x) + 5)^2 - (2*exp(x)*(exp(x) + x*exp(x) + 5))/(exp(x) + 5)^3